# Prime and Co-prime Integer Matrices

### v2

@@ I Introduction
It is well-known that prime integers are the most important subjects in mathematics and have played vital roles in applications, such as cryptography, digital communications, and signal processing etc. It plays the key role in the Chinese remainder theorem (CRT) that is to reconstruct a large integer from its remainders modulo several small moduli and has many applications in cryptography, digital communications and signal processing as well [1]. For a given set of moduli, the largest integer that can be uniquely reconstructed from its remainders is smaller than the least common multiple (lcm) of all the moduli and the lcm achieves the maximal if all the moduli are pairwise co-prime, when the moduli sizes are upper bounded. This means that it is critical to have a large size set of pairwise co-prime integers.
@@
In this paper, we first recall prime integer matrices and some of their basic properties presented in [2]. We then investigate co-prime and pairwise co-prime integer matrices. We show that, similar to prime integers, two different prime integer matrices are co-prime. We then present the Hermite normal forms [14] for prime integer matrices, and show that two prime integer matrices with the same determinant absolute value but different Hermite normal forms are co-prime. Thus, it is easy to construct families of pairwise co-prime integer matrices by simply contructing families of prime integer matrices with different determinant absolute values that are prime and/or with the same determinant absolute value that is prime but different Hermite normal forms. This characterizes all pairwise co-prime integer matrices that are all prime. It is interesting to note that none of the pairwise co-prime integer matrices in the construction in [17] is prime when the dimension is above $1$. Also note that a different kind of primality for matrices was studied in [12, 13] for a given set of matrices.
Since Gaussian integers are special $2\times 2$ integer matrices, interestingly, Gaussian primes and prime $2\times 2$ integer matrices are not exactly the same. When both real and imaginary parts of Gaussian integers are not 00, Gaussian primes and prime $2\times 2$ integer matrices are the same. Otherwise, Gaussian primes may not be prime $2\times 2$ integer matrices. However, two Gaussian integers are co-prime if and only if their equivalent $2\times 2$ integer matrices are co-prime. It is well-known that any Gaussian integer has a unique Gaussian prime factorization [3]. In contrast, an integer matrix can be unqiuely factorized to a product of prime integer matrices when the order of the prime integer matrices in the product is not considered. Note that the order of primes in a product is not a problem if the primes are communtative, such as Gaussian primes (or their equivalent $2\times 2$ integer matrices).
@@
II Definitions and Properties of Prime Integer Matrcies
@@
For the consistency, $1$ is considered as a conventional prime integer, which corresponds to a unimodular matrix treated as a prime matrix in this paper. It is clear that if $A$ is prime, then $AU$ and $UA$ are both prime for any unimodular matrix $U$. Since the non-commutativity of matrices and we only consider one side primality, i.e., the left primality. For this reason, in this paper, a matrix $A$ and its right multiples $AU$ for unimodular matrices $U$ are the same (or in an equivalent class). In other words, matrix $A$ and matrix $AU$ for any unimodular matrix $U$ are indistinguishable. Clearly, matrices with different determinant absolute values are different. In this sense, all the lcrm of a set of matrices $M_{1},M_{2},\dots,M_{L}$ are the same, i.e., the lcrm of a given set of matrices is unique.
@@
We next provide a complete answer to this question by using Hermite normal form of integer matrices (Chapter 14 of [2] and Chapter 14 of [14]) where is says that any matrix has a unique Hermite normal form. Combined with the primality of a matrix with the result in Corollary 1, we have the following proposition.
@@
For the MD-CRT, similar to the conventional CRT, the range of the uniquely determineable integer vectors from their vector remainders depends on an lcrm of the matrix moduli. Similar to the conventional integer case, we have the following proposition.
@@
Interestingly, algebraic integers in a quadraic algebraic number field with a minimal polynomial $r_{2}x^{2}+r_{1}x+r_{0}$ can be equivalently represented by $2\times 2$ matrices of the form in (3) with $\alpha=r_{0}/r_{2}$ and $\beta=r_{1}/r_{2}$, where $b$ can be 00, [10, 11]. Since algebraic integers are all complex numbers and therefore naturally commutative. When $r_{2}=1,r_{1}=0,r_{0}=1$, the above algebraic integers become Gaussian integers that will be especially studied in next section.
@@ III Connection with Gaussian Integers and Gaussian Primes
In other words, matrix primality implies Gaussian primality, and Gaussian primality may not imply matrix primality. One reason might be as follows. Since in terms of the matrix primality, a matrix can be equivalently converted to a diagonal matrix of the form in Corollary 1 by using the Smith form decomposition that may not always preserve the matrix representation structure (4) for a Gaussian integer in the decompsition steps. This means that the Smith form in Corollary 1 for a prime matrix may not apply to a Gaussian prime, i.e., a Gaussian prime does not have to have the Smith form in Corollary 1.
We next want to see the co-primality of two matrices and two Gaussian integers. First two Gaussian integers are called co-prime if they do not have any non-unit common factor. Although the primalities of a Gaussian integer and its matrix representation are not equivalent, we will show below that the co-primalities of two Gaussian integers and their matrix representations are equivalent.
@@
Two Gaussian integers are co-prime if and only if their matrix representations are co-prime in the integer matrix sense.
Proof. We first note that for both integer matrices and Gaussian integers, Bezout’s identity holds. Let $z_{i}$ have matrix representations $Z_{i}$ in (4) for $i=1,2$. From Bezout’s identity, $z_{1}$ and $z_{2}$ are co-prime if and only if there exist Gaussian integers $y_{1}$ and $y_{2}$ such that $z_{1}y_{1}+z_{2}y_{2}=q$ for a unit $q$. Let $Y_{1},Y_{2},Q$ be the matrix representations of $y_{1},y_{2},q$ in (4), respectively. Then, we have $Z_{1}Y_{1}+Z_{2}Y_{2}=Q$ and vice versa. Since $q$ is a unit if and only if $Q$ is unimodular, $z_{1}y_{1}+z_{2}y_{2}=q$ for a unit $q$ if and only if $Z_{1}Y_{1}+Z_{2}Y_{2}=Q$ for a unimodular matrix $Q$ if and only if $Z_{1}$ and $Z_{2}$ are co-prime by Bezout’s indentity for integer matrices [9]. q.e.d.

### v3

@@ I Introduction
It is well-known that prime integers are the most important subjects in mathematics and have played vital roles in applications, such as cryptography, digital communications, and signal processing etc. It plays a key role in the Chinese remainder theorem (CRT) that is to reconstruct a large integer from its remainders modulo several small moduli and has many applications in cryptography, digital communications and signal processing as well [1]. For a given set of moduli, the largest integer that can be uniquely reconstructed from its remainders is smaller than the least common multiple (lcm) of all the moduli and the lcm achieves the maximal if all the moduli are pairwise co-prime, when the moduli sizes are upper bounded. This means that it is critical to have a large size set of pairwise co-prime integers.
@@
In this paper, we first recall prime integer matrices and some of their basic properties presented in [2]. We then investigate co-prime and pairwise co-prime integer matrices. We show that, similar to prime integers, two different prime integer matrices are co-prime. We then present the Hermite normal forms [14] for prime integer matrices, and show that two prime integer matrices with the same determinant absolute value but different Hermite normal forms are co-prime. Thus, it is easy to construct families of pairwise co-prime integer matrices by simply constructing families of prime integer matrices with different determinant absolute values that are prime and/or with the same determinant absolute value that is prime but different Hermite normal forms. This characterizes all pairwise co-prime integer matrices that are all prime. It is interesting to note that none of the pairwise co-prime integer matrices in the construction in [17] is prime when the dimension is above $1$. Also note that a different kind of primality for matrices was studied in [12, 13] for a given set of matrices.
Since Gaussian integers are special $2\times 2$ integer matrices, interestingly, Gaussian primes and prime $2\times 2$ integer matrices are not exactly the same. When both real and imaginary parts of Gaussian integers are not 00, Gaussian primes and prime $2\times 2$ integer matrices are the same. Otherwise, Gaussian primes may not be prime $2\times 2$ integer matrices. However, two Gaussian integers are co-prime if and only if their equivalent $2\times 2$ integer matrices are co-prime. It is well-known that any Gaussian integer has a unique Gaussian prime factorization [3]. In contrast, an integer matrix can be uniquely factorized to a product of prime integer matrices when the order of the prime integer matrices in the product is not considered. Note that the order of primes in a product is not a problem if the primes are commutative, such as Gaussian primes (or their equivalent $2\times 2$ integer matrices).
@@
II Definitions and Properties of Prime Integer Matrices
@@
For the consistency, $1$ is considered as a conventional prime integer, which corresponds to a unimodular matrix treated as a prime matrix in this paper. It is clear that if $A$ is prime, then $AU$ and $UA$ are both prime for any unimodular matrix $U$. Since the non-commutativity of matrices and we only consider one side primality, i.e., the left primality. For this reason, in this paper, a matrix $A$ and its right multiples $AU$ for unimodular matrices $U$ are counted the same (or associates [2]) for simplicity. In other words, matrix $A$ and matrix $AU$ for any unimodular matrix $U$ are indistinguishable. Clearly, matrices with different determinant absolute values are different. In this sense, all the lcrm of a set of matrices $M_{1},M_{2},\dots,M_{L}$ are the same, i.e., the lcrm of a given set of matrices is unique.
@@
We next provide a complete answer to this question by using Hermite normal form of integer matrices (Chapter 14 of [2] and Chapter 14 of [14]) where it says that any matrix has a unique Hermite normal form. Combined with the primality of a matrix with the result in Corollary 1, we have the following proposition.
@@
For the MD-CRT, similar to the conventional CRT, the range of the uniquely determinable integer vectors from their vector remainders depends on an lcrm of the matrix moduli. Similar to the conventional integer case, we have the following proposition.
@@
Interestingly, algebraic integers in a quadratic algebraic number field with a minimal polynomial $r_{2}x^{2}+r_{1}x+r_{0}$ can be equivalently represented by $2\times 2$ matrices of the form in (3) with $\alpha=r_{0}/r_{2}$ and $\beta=r_{1}/r_{2}$, where $b$ can be 00, [10, 11]. Since algebraic integers are all complex numbers and therefore naturally commutative. When $r_{2}=1,r_{1}=0,r_{0}=1$, the above algebraic integers become Gaussian integers that will be especially studied in next section.
@@ III Connection with Gaussian Integers and Gaussian Primes
In other words, matrix primality implies Gaussian primality, and Gaussian primality may not imply matrix primality. One reason might be as follows. Since in terms of the matrix primality, a matrix can be equivalently converted to a diagonal matrix of the form in Corollary 1 by using the Smith form decomposition that may not always preserve the matrix representation structure (4) for a Gaussian integer in the decomposition steps. This means that the Smith form in Corollary 1 for a prime matrix may not apply to a Gaussian prime, i.e., a Gaussian prime does not have to have the Smith form in Corollary 1.
We next want to see the co-primality of two matrices and two Gaussian integers. First two Gaussian integers are called co-prime if they do not have any non-unit common factor. Although the primalities of a Gaussian integer and its matrix representation are not equivalent, the co-primalities of two Gaussian integers and their matrix representations are equivalent [10].
@@
[10] Two Gaussian integers are co-prime if and only if their matrix representations are co-prime in the integer matrix sense.
From the above results, it is not hard to see that two Gaussian primes with different norms are co-prime, similar to the conventional prime integers. For two Gaussian primes with the same norm, their equivalent matrix representations have the same determinant absolute value. In this case, from the $2\times 2$ integer matrix point of view, two different prime integer matrices with the same determinant absolute value may be co-prime if their Hermite normal forms are different from Prop. 4. If they only differ by a unit, then they are the same and thus not co-prime, and their equivalent matrix representations have the same Hermite normal form. If they do not differ by a unit, they have to be co-prime. Then, their equivalent matrix representations have different Hermite normal forms. An exmaple is $4+5j$ and $4-5j$ that have the same norm but different Hermite normal forms and thus co-prime, and do not differ by a unit.